# Low-Cost Architecture and Efficient Pattern Synthesis for Polarimetric Phased Array Based on Polarization Coding Reconfigurable Elements

### v2

@@ III-C Dual-polarized Pattern Synthesis
For the measurement applications of the STSR mode, a pair of orthogonally polarized beams is required simultaneously. Therefore, all the elements are split into two groups to synthesize the beams. This can be represented by the waveform coding vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, where $x_{1},x_{2}\in\{0,1\}$ and $x_{i,j}=1$ means the $i$-th element belongs to the $j$-th beam. Take the beam $1$ as an example, Equation (10) can be transferred to:
@@
In this case, the synthesis of dual-polarized beam patterns is decided by variables of the beam coding vectors $\boldsymbol{x}_{1h}$, $\boldsymbol{x}_{1v}$, $\boldsymbol{x}_{2h}$, and $\boldsymbol{x}_{2v}$, which are defined by the Hadamard product of the polarization and waveform coding vectors.
According to [40], the calibration methods can conveniently mitigate the cross-polarization components of these dual-polarized beams. Therefore, it is reasonable to assume that no cross-polarized elements are involved in each beam, which can be expressed as:
$\left\{{\begin{array}[]{l}{\boldsymbol{x}_{1}=\boldsymbol{x}_{h}}\\ {\boldsymbol{x}_{2}=\boldsymbol{x}_{v}}\end{array}}\right$. (23)
Since Equation (23) holds, the variables $\boldsymbol{x}_{1v}$ and $\boldsymbol{x}_{2h}$ are equal to zero, from which can infer that:
$\boldsymbol{x}_{1h}+\boldsymbol{x}_{2v}=\boldsymbol{1}$ (24)
where $\text{sum}(\boldsymbol{x}_{1h})=\text{sum}(\boldsymbol{x}_{2v})=N/2$ to ensure good match of the co-polar beams.
A Monte Carlo simulation given by Fig. 5 has been carried out according to Equation (24) to investigate the PSL distribution under the dual-polarized pattern synthesis model. The beampoint is set at $(120^{\circ},15^{\circ})$, with the PSL taking the maximum of the H and V beams. The red dot line marks the value calculated under the CPPA using Equation (6). Unlike the phenomenon revealed by Fig. 4, this histogram shows a more dispersed distribution, with the worst PSL 5 dB higher than the CPPA. In addition, the fitting curve marked by the red solid line indicates that only 0.92% of the configurations have PSLs that are no worse than that of the CPPA.
@@
Therefore, an optimized model with a central symmetry constraint is introduced to derive low PSL results. This model can be solved using binary genetic algorithm (BGA) methods and has been discussed in detail in previous work by the authors[35]. The central symmetry constraint means that two configurations forming the H and V beams are centrally symmetric shown in Fig. 6. In this case, the array factors produced by these configurations will have the same power patterns, which have been theoretically demonstrated in [35].
@@
$\boldsymbol{X}_{2v}=\boldsymbol{1}-\text{rot180}(\boldsymbol{X}_{1h})$ (25)
@@
From the perspective of matrix operations, Equation (25) can be further represented in vector form using the code vectors $\boldsymbol{x}_{1h}$ and $\boldsymbol{x}_{2v}$:
@@
The pattern synthesis problem can be then formulated as a minimax problem for minimizing the PSL of $f_{1h}\left(\theta,\phi\right)$ and $f_{2v}\left(\theta,\phi\right)$ by determining the coding of elements in the binary vector $\boldsymbol{x}_{u}$, which is given by
@@
This nonlinear integer programming problem tends to be solved by stochastic optimization algorithms such as BGA, binary PSO (BPSO), and binary differential evolution (BDE). In this paper, the BGA algorithm is adopted. The whole procedure of the synthesis method is shown in Algorithm 2 and explained step by step.
@@
2: Optimize the variable $\boldsymbol{x}_{u}$ using the BGA algorithm according to the model given by Equation (27). Stop the process if it reaches the termination condition.
3: Calculate the optimized configuration results according to Equation (26).
4: Calibrate the dual-polarized patterns according to the method mentioned in [40].
III-D Simulation Results
Input data from the HFSS model of a $16\times 16$-unit dual-polarized microstrip array antenna is adopted to demonstrate the performance of the proposed architecture and methods. The simulations are carried out for vertically, arbitrarily, and dual-polarized patterns concerning various radar applications.
III-D1 Vertically Polarized Pattern
This simulation considers a V polarization pattern with the beam direction at $(120^{\circ},15^{\circ})$. These beams with pure V or H polarization are standard in polarization-measuring applications of ATSR modes. According to the initial conditions and the method given in Algorithm 1, the results of the element polarization configuration are shown in Fig. 7. In this configuration, 17 out of 256 elements, which account for 6.64%, are excited by the H-pol ports to eliminate the cross-polarization components of the synthesized beam. Their positions are generated based on the proposed method.
Fig. 8 presents the 3-D normalized power patterns of different array architectures or methods under the same data input and initial conditions. The calibration method aimed at the ATSR mode is adopted from [40]. Compared to the CPPA, PCRPA exhibits a noticeable improvement in XP component mitigation, with the value of the cross-polarization level XPL decreasing from -22.93 dB to -59.68 dB, as shown in Fig. 8b and 8f. This is comparable to the effects of calibration, as shown in Fig. 8d, where the XPL is -65.31 dB. Meanwhile, the CoP pattern of PCRPA indicates only an even better level in the PSL compared to conventional PPA, as shown in Fig. 8a, 8c, and 8e.
Figure 7: H-pol and V-pol element distribution of a $16\times 16$-unit planar array for the required polarization state $(90^{\circ},0^{\circ})$ and beam direction $(120^{\circ},15^{\circ})$. The blue and red squares represent H-pol and V-pol elements, respectively.
Figure 8: The 3-D normalized power patterns of the simulated $16\times 16$-unit microstrip antenna array, where V polarization is the co-polarization and the beam direction is $(120^{\circ},15^{\circ})$. The red dashed line helps indicate the pointing direction of the array beam. (a) The CoP pattern before calibration of the CPPA. (b) The XP pattern before calibration of the CPPA. (c) The CoP pattern after calibration of the CPPA. (d) The XP pattern after calibration of the CPPA. (e) The CoP pattern of the proposed PCRPA. (f) The XP pattern of the proposed PCRPA.
TABLE II: Comparison of the key parameters of the vertically polarized pattern synthesis
Group CPPA, uncalibrated CPPA, calibrated PCRPA
XPL/dB -22.93 -65.31 -59.68
PSL/dB -12.75 -12.77 -13.39
Pt/dB 0 -0.01 -0.55
Dir/dBi 28.37 28.37 28.15
To clarify in detail, Table III presents the specific parameter values of the three patterns. Here, $P_{t}$ represents the transmission power at the beam direction normalized with the value of the uncalibrated beam under CPPA; $D$ denotes the directivity of the array pattern. As indicated by the data, since some elements are synthesized to suppress XP components, the integrating sidelobes of both the CoP and XP patterns of PCRPA have risen, leading to a loss of power and directivity. The results show that $P_{t}$ and $D$ of PCRPA are reduced by 0.55 dB and 0.22 dB, respectively.
III-D2 Arbitrarily Polarized Pattern
This simulation considers an arbitrarily polarized pattern with the same beam direction as in the above section. Such requirements tend to appear in many anti-interference applications, where the beam polarization is orthogonal to the interference signal polarization to suppress its received power. Here, the desired polarization state is set to $(30^{\circ},60^{\circ})$. According to Equation (13), the 2-D normalized power patterns of the CPPA are given by Fig. 9. The XPL of this beam is at a relatively high level of -21.25 dB, which may affect the anti-interference processing.
Figure 9: The azimuth cut plane of the 2-D normalized power pattern of the arbitrarily polarized pattern synthesis under the CPPA architecture, with the array size of $16\times 16$, co-polarization state at $(30^{\circ},60^{\circ})$ and the beam direction at $(120^{\circ},15^{\circ})$.
Fig. 10 presents three groups of PCRPA results obtained by different pattern synthesis methods, including the proposed, random, and BGA methods. Table III lists the key parameters of these results and the CPPA in detail. Here, $\text{XPL}_{\text{m}}$ denotes the normalized peak level of the XP pattern across the scan principal planes. The difference between these three methods lies in Step 4 to Step 6 of the algorithm given in subsection III-B. The random method obtains the array configuration by performing the random number generator only once, where the XPL and PSL constraints are ineffective. The BGA method obtains the array configuration by optimizing the coding vectors $\boldsymbol{x}_{h}$ and $\boldsymbol{x}_{v}$ using a binary genetic algorithm. Computations were performed on a PC with an Intel Core i7-10510U central processing unit (CPU) at 2.30 GHz and 16 GB RAM.
Figure 10: Comparison of the arbitrarily polarized pattern synthesis results for the proposed architecture using different methods. (a) Elevation plane using the proposed method. (b) Elevation plane using the random method. (c) Elevation plane using the GA method. (d) Azimuth plane using the proposed method. (e) Azimuth plane using the random method. (f) Azimuth plane using the GA method. (g) Array configuration using the proposed method. (h) Array configuration using the random method. (i) Array configuration using the GA method.
TABLE III: Comparison of key parameters of the arbitrarily polarized pattern synthesis
Group CPPA PCRPA
proposed random BGA
XPL/dB -24.43 -49.13 -49.13 -49.05
$\text{XPL}_{\text{m}}$/dB -24.15 -20.29 -8.51 -20.61
PSL/dB -12.75 -13.38 -12.63 -13.93
Run time/s \ 5.19 \ 1206.95
Considering the comparison of XPL and PSL between CPPA and PCRPA, the simulation results here are consistent with the V polarization simulations. The advantage of PCRPA’s cross-polarization suppression still exists, and so does the decrease of the peak sidelobe, which is at a value of -13.38 dB. Besides, among the 3 groups of PCRPA results, the XP patterns of the random method show an excessively high level at -8.71 dB within the main lobe area, given in Fig. 10b and 10e. This excessively high sidelobe of XP patterns is eliminated in the other 2 groups due to the optimization of element configurations, as shown in Fig. 10a, 10c, 10d, and 10f. Furthermore, the proposed method and the BGA method’s pattern synthesis results indicate comparable performance, with the proposed method’s PSL being 0.55 dB higher. However, its run time is also significantly shorter, only 5.19 seconds compared with 1206.95 seconds using the BGA method according to Table III, which is an obvious efficiency advantage. The simulation results presented above show that the proposed method can achieve a compromise between performance and efficiency in the PCRPA architecture.
III-D3 Dual-polarized Patterns
Figure 11: H-pol and V-pol element distribution of a $16\times 16$-unit planar array for dual-polarized beams with beam direction $(120^{\circ},15^{\circ})$. The blue squares and red diamonds represent H-pol and V-pol elements, respectively.
Figure 12: Comparison of the dual-polarized pattern synthesis results for the conventional and proposed architecture using different methods. (a) Elevation plane of the CPPA. (b) Elevation plane of the PCRPA using the proposed method. (c) Elevation plane of the PCRPA using the random method. (d) Azimuth plane of the CPPA. (e) Azimuth plane of the PCRPA using the proposed method. (f) Azimuth plane of the PCRPA using the random method.
TABLE IV: Comparison of different groups
Group CPPA PCRPA
proposed random
XPL/dB H -56.89 -56.76 -56.86
V -57.80 -57.87 -57.84
PSL/dB H -12.74 -12.73 -11.70
V -12.77 -12.76 -8.95
$P_{t}$/dB H 0 -6.02 -6.02
V -0.02 -6.04 -6.04
DD/dBi H 28.37 26.12 25.86
V 28.37 26.12 26.07
ME/dB -19.57 -19.59 10.36
This simulation considers a pair of H and V polarization beams with the same beam direction as above, which are commonly required in polarization measuring applications of the STSR mode. According to the proposed synthesis method given by Equation (27) and Algorithm 2, the entire array is split into two groups, each fed with the corresponding waveform. To quantitatively evaluate the matching performance of the main lobes in dual-polarized beams, a new parameter called matching error (ME) is introduced, defined as[41]:
[…]
where $f_{1h,k}$ and $f_{2v_{k}}$ represent the $k$-th sampling points within the main lobe of patterns $f_{1h}(\theta,\phi)$ and $f_{2v}(\theta,\phi)$.
Fig. 12 compares three groups of dual polarization patterns obtained under different configurations: the CPPA, the optimized PCRPA, and a bad case of PCRPA. All the patterns have been processed by polarization calibration methods. Fig. 11 shows the configuration of the optimized PCRPA. Markers of squares and diamonds represent the H and V polarization excited elements to distinguish the coding of waveforms. The bad case configuration is obtained without the central symmetry constraint proposed in this paper, simply using a random number generator instead. The optimized PCRPA performs well on the matched co-polar beams shown in Fig. 12b and 12e. Its ME value is very close to that of conventional PPA and 29.96 dB better than that of the bad case, which verifies the effectiveness of the constraint. The same phenomenon also applies to the PSL parameter detailed in Table IV. In addition to the above advantages, Table IV shows that the PCRPA has a significant loss problem on power and directivity at around 6 dB and 2.2 dB, respectively, compared to the CPPA. This may limit the maximum radar range under STSR mode.
IV Parameter Sensitivity Analysis
Figure 13: Pattern performance of the proposed and conventional polarimetric phased arrays under different desired polarization states. (a) PSL and XPL under different $\gamma$. (b) Pt and Dir under different $\gamma$. (c) PSL and XPL under different $\eta$. (d) Pt and Dir under different $\eta$.

### v3

@@ III-C Dual-polarized Pattern Synthesis
For the measurement applications of the STSR mode, a pair of orthogonally polarized beams is required simultaneously. Therefore, all the elements are split into two groups to synthesize the beams. This can be represented by the subarray coding vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, where $x_{1},x_{2}\in\{0,1\}$ and $x_{i,j}=1$ means the $i$-th element belongs to the $j$-th beam. Take the beam $1$ as an example, Equation (11) can be transferred to:
@@
The model given in Equation (23) involves nonlinear vector dot products, making direct solution computationally complex. A divide-and-conquer approach is therefore adopted. The problem is decomposed into two sequential steps: first, the subarray coding vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{1}$, which determine the co-polar pattern performance for dual beams, are synthesized. Subsequently, the polarization coding vectors $\boldsymbol{x}_{h}$ and $\boldsymbol{x}_{v}$ are optimized using Algorithm 1 described in the previous subsection for arbitrary polarization pattern synthesis, with detailed steps provided later.
A Monte Carlo simulation given by Fig. 5 has been carried out to investigate the PSL distribution under the dual-polarized pattern synthesis model. According to the divide-and-conquer approach, assume that $\boldsymbol{x}_{1}=\boldsymbol{x}_{h}=\boldsymbol{x}_{1h}$ and $\boldsymbol{x}_{2}=\boldsymbol{x}_{v}=\boldsymbol{x}_{2v}$. The beampoint is set at $(120^{\circ},15^{\circ})$, with the PSL taking the maximum of the H and V beams. The red dot line marks the value calculated under the CPPA using Equation (7). Unlike the phenomenon revealed by Fig. 4, this histogram shows a more dispersed distribution, with the worst PSL 5 dB higher than the CPPA. In addition, the fitting curve marked by the red solid line indicates that only 0.92% of the configurations have PSLs that are no worse than that of the CPPA.
@@
Therefore, an optimized model with a central symmetry constraint is introduced to derive low PSL results. This model can be solved using binary genetic algorithm (BGA) methods and has been discussed in detail in previous work by the authors[wang2024low-cost]. The central symmetry constraint means that two configurations forming the H and V beams are centrally symmetric shown in Fig. 6. In this case, the array factors produced by these configurations will have the same power patterns, which have been theoretically demonstrated in [wang2024low-cost].
@@
$\boldsymbol{X}_{2}=\boldsymbol{1}-\text{rot180}(\boldsymbol{X}_{1})$ (24)
@@
From the perspective of matrix operations, Equation (24) can be further represented in vector form using the subarray code vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$:
@@
The pattern synthesis problem can be then formulated as a minimax problem for minimizing the PSL of $f_{1}\left(\theta,\phi\right)$ and $f_{2}\left(\theta,\phi\right)$ by determining the coding of elements in the binary vector $\boldsymbol{x}_{u}$, which is given by
@@
This nonlinear integer programming problem is solved using a stochastic optimization algorithm, specifically the BGA adopted in this work.
For the $i$-th subarray coding vector $\mathbf{x}_{i}\in\{0,1\}^{n}$, a transformation matrix $\mathbf{L}_{i}\in\{0,1\}^{n_{i}\times n}$ is constructed, where $n_{i}=\|\mathbf{x}_{i}\|_{1}$ denotes the number of active elements in this subarray. Let $\mathbf{Q}=\{q_{1},q_{2},\ldots,q_{n_{i}}\}$ be the ordered set of indices where the components of $\mathbf{x}_{i}$ are equal to 1. The transformation matrix $\mathbf{L}_{i}$ is then defined element-wise by:
[…]
for row index $k=1,2,\ldots,n_{i}$ and column index $l=1,2,\ldots,n$. This matrix maps from the full array’s indices to the indices of the given subarray code vector.
Since the dual-beam synthesis problem is decomposed into two independent single-beam synthesis subproblems, for the $i$-th beam ($i=1,2$), the input data can be obtained by mapping the original AEP data $\mathbf{g}_{AB}$ via the transformation matrix $\mathbf{L}_{i}$:
$\tilde{\mathbf{g}}_{AB}^{(i)}(\theta,\phi)=\mathbf{L}_{i}\mathbf{g}_{AB}(\theta,\phi)$ (28)
Given the constraints including the beam direction, polarization state, and sidelobe level, Algorithm 1 is performed to optimize $\tilde{\mathbf{g}}_{AB}^{(i)}(\theta,\phi)$, yielding the local solutions $\mathbf{x}_{h}^{i}$ and $\mathbf{x}_{v}^{i}$ corresponding to this subarray.
Upon completion of the optimization for all beams, the local solutions are combined into the final code vectors for the full array via the transpose of the transformation matrices:
$\left\{\begin{aligned} \mathbf{x}_{ih}=\mathbf{L}_{i}^{\text{T}}\mathbf{x}_{h}^{i}\\ \mathbf{x}_{iv}=\mathbf{L}_{i}^{\text{T}}\mathbf{x}_{v}^{i}\end{aligned}\right$. (29)
The vectors $\mathbf{x}_{ih}$ and $\mathbf{x}_{iv}$ constitute the final output of the dual-beam pattern synthesis process.
Consistent with the approach in Algorithm 1, the synthesis procedure given by Algorithm 2 is also founded on the AEP model. By utilizing full-wave simulation or measured data for array pattern calculation, this method inherently accounts for mutual coupling effects. The algorithm steps are explained as follows.
@@
2: Optimize the variable $\boldsymbol{x}_{u}$ using the BGA algorithm according to the model given by Equation (26). Stop the process if it reaches the termination condition and obtain the subarray code vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$.
3: Perform Algorithm 1 under $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ respectively according to Equation (27)-(29). Return the synthesis results.
IV Simulation And Analysis
This section presents simulation examples to validate the proposed PCRPA architecture and its pattern synthesis methods. The simulation begins with acquiring the AEP data of all ports from a dual-polarized array model through full-wave analysis using the high-frequency structure simulator (HFSS). A dual-polarized microstrip patch antenna, operating at a center frequency of 10 GHz, was employed as the array element. The square patch has a side length of 8.3 mm and is fed by coaxial probes. The array lies in the yoz-plane, with all elements arranged in a $16\times 16$ uniform rectangular lattice with an element spacing of 15 mm (i.e., half-wavelength at 10 GHz). Since this section focuses on validating the proposed architecture and methodology, the electromagnetic simulation was configured for a single frequency of 10 GHz. These data are then used as input to the synthesis model, where the binary coding vectors are optimized to emulate the switching functionalities of the architecture in Fig. 2.
A set of three key metrics is employed to evaluate the pattern performance: PSL, XPL, and directivity (Dir). PSL and Dir characterize the concentration of co-polarized radiation energy, while XPL quantifies the polarization purity at the beam direction. These metrics are consequently adopted for clear comparison in following simulations and experiments.
IV-A Arbitrarily Polarized Pattern
This set of simulations evaluates the performance of Algorithm 1 for synthesizing arbitrarily-polarized beams steered to $(120^{\circ},15^{\circ})$. The synthesis results of array configurations and beam patterns for two polarization states, $(90^{\circ},0^{\circ})$ and $(45^{\circ},0^{\circ})$, are presented in Fig. 7 as examples.
Figure 7: Simulation results for arbitrarily polarized pattern synthesis. (a) Array configuration (V-polarization). (b) Array configuration ($45^{\circ}$ linear polarization). (c) CoP pattern of V-polarized beam. (d) XP pattern of V-polarized beam. (e) CoP pattern of $45^{\circ}$ linearly-polarized beam. (f) XP pattern of $45^{\circ}$ linearly-polarized beam.
In Fig. 7a and 7b, orange and blue colors correspond to the H and V port excitations, respectively. The results show that the PCRPA uses 17 H ports and 239 V ports to synthesize the V-polarized beam, whereas for the $45^{\circ}$ linearly polarized beam, the numbers change to 159 and 97. Both configurations form stable main lobes at the desired direction with controlled PSL and obvious XPL suppression. However, the XP pattern of the $45^{\circ}$ linearly polarized beam exhibits a higher level across the entire space, indicating a sensitivity of the synthesis performance to the polarization state, which will be further discussed later.
Fig. 8 further compares the elevation plane cuts of the PCRPA patterns from Fig. 7 with those from the CPPA and [zhou2024], with the corresponding quantitative metrics summarized in TABLE III. The comparative results show that for V-polarized synthesis, both [zhou2024] and the PCRPA maintain a PSL comparable to the CPPA, albeit with directivity loss at approximately 0.2 dB due to elevated integrated sidelobes. In terms of cross-polarization, the architecture in [zhou2024] exhibits an XPL similar to the CPPA, with its XP pattern generally elevated across the entire spatial domain. In contrast, the PCRPA achieves significant suppression of both the XPL and the cross-polarization within the main lobe. For $45^{\circ}$ linearly-polarized synthesis, the PCRPA maintains the same advantages observed in the V-polarized case; however, its level of XP pattern outside the main lobe is further increased. It is noteworthy that the architecture proposed in [zhou2024] cannot form polarization states other than H and V under uniform weighting; therefore, its performance for $45^{\circ}$ linear polarization could not be compared under identical conditions.
Figure 8: Normalized 2-D power patterns (elevation plane cut) for dual-polarized beams. (a) CoP patterns of V-polarized beams. (b) XP patterns of V-polarized beams. (c) CoP patterns of $45^{\circ}$ linearly-polarized beams. (d) XP patterns of $45^{\circ}$ linearly-polarized beams.
TABLE III: Comparison of key metrics for the arbitrarily polarized pattern synthesis
CPPA Ref. [zhou2024] PCRPA
$(90^{\circ},0^{\circ})$ XPL (dB) -22.93 -22.89 -59.68
PSL (dB) -12.75 -12.70 -13.39
Dir (dBi) 28.37 28.14 28.15
$(45^{\circ},0^{\circ})$ XPL (dB) -24.43 -49.13
PSL (dB) -12.75 -13.38
Dir (dBi) 28.37 28.00
Amplitude Weight Uniform Uniform Uniform
IV-B Dual-polarized Patterns
This set of simulations evaluates the performance of the proposed dual-polarized beam synthesis method with beams steered to $(120^{\circ},15^{\circ})$. The sythesis results obtained are presented in Fig. 9. Diamonds and squares denote the subarray excitations for the H- and V-polarized beams, respectively; orange and blue colors correspond to the H and V port excitations of the elements. As shown, each beam is synthesized using half elements of the total array. The H-polarized beam is excited by 120 H ports and 8 V ports, whereas the V-polarized beam is excited by 120 V ports and 8 H ports. Fig 9 also presents the CoP and XP patterns for both beams. It can be observed that both CoP patterns form stable main lobes at the desired beam direction, with the PSL below -13 dB, and the XPL suppressed to a low level. Conversely, the integrated sidelobe levels of the CoP patterns are higher than those in the single-beam case, leading to a measurable loss in directivity.
Figure 9: Simulation results for dual-polarized beam synthesis. (a) Array configuration. (b) CoP pattern of H-polarized beam. (c) XP pattern of H-polarized beam. (d) CoP pattern of the V-polarized beam. (e) XP pattern of the V-polarized beam.
Fig. 10 presents the elevation plane cuts of the CoP and XP patterns for the PCRPA dual-beam array from Fig. 9, comparing them with the dual-polarized beams from the CPPA and a randomly coded control group. The corresponding performance metrics are compared in TABLE IV. The control group was configured such that its subarray coding vectors, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, were randomly generated while maintaining an equal number of excited elements, with $\boldsymbol{x}_{1}=\boldsymbol{x}_{1h}$ and $\boldsymbol{x}_{2}=\boldsymbol{x}_{2v}$. This configuration excludes the geometric constraints, element position optimization, and polarization port allocation strategies in Algorithm 2. The results demonstrate that the PCRPA optimized by the proposed method maintains comparable PSL while achieving better XPL over the other. The XPL improved from -22.46/-22.93 dB to -41.55/-45.78 dB. Furthermore, compared to the random control group, the proposed algorithm yields well-matched CoP patterns for the H- and V-beams, as evidenced by their nearly identical main lobe widths in Fig. 10. However, a directivity loss of approximately 2.50 dB is observed in the PCRPA architecture, which is an inherent trade-off due to the use of only a subset of elements for each beam.
Figure 10: Normalized 2-D power patterns (elevation plane cut) for dual-polarized beams. (a) CoP patterns. (b) XP patterns.
TABLE IV: Comparison of key metrics for the dual polarized pattern synthesis
CPPA PCRPA
random proposed
XPL (dB) H -22.46 -22.46 -41.55
V -22.93 -22.94 -45.78
PSL (dB) H -12.73 -11.68 -13.31
V -12.75 -8.91 -13.48
Dir (dBi) H 28.37 25.84 25.87
V 28.37 26.05 25.86
IV-C Parameter Sensitivity Analysis
Figure 11: Comparison of pattern performance between the conventional and proposed architecture under varying polarization states. (a)-(d) show XPL, PSL, Dir, and $P_{t}$ as functions of $\gamma$; (e)-(h) present the same metrics as functions of $\eta$.